OpenStudy (anonymous):

Find the slope of the line passing through each pair of points & explain how you solved. (-6, 1) and (3, -2)

OpenStudy (anonymous):

\[m=\frac{-2-1}{3-(-6)}=\frac{-3}{3+6}\]

OpenStudy (anonymous):

(1-(-2))/(-6-3)=-1/3

OpenStudy (anonymous):

\[\frac{-3}{9}=\frac{-1}{3}\]

OpenStudy (anonymous):

and do you use \[y^{2} - y^{1} \] over x^2 - x^1?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay. is the slope -1/3?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank for medal